# Can LLMs handle WebShell detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v1

@@ Abstract
WebShell attacks, in which malicious scripts are injected into web servers, are a major cybersecurity threat. Traditional machine learning and deep learning methods are hampered by issues such as the need for extensive training data, catastrophic forgetting, and poor generalization. Recently, Large Language Models (LLMs) have gained attention for code-related tasks, but their potential in WebShell detection remains underexplored. In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain. Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles. Our results show that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off. However, all models lag behind previous State-Of-The-Art (SOTA) methods. With BFAD, the performance of all LLMs improved, with an average F1 score increase of 13.82%. Larger models such as GPT-4, LLaMA 3.1 70B, and Qwen 2.5 14B outperform SOTA methods, while smaller models such as Qwen 2.5 3B achieve performance competitive with traditional methods. This work is the first to explore the feasibility and limitations of LLMs for WebShell detection, and provides solutions to address the challenges in this task.
@@ 1 Introduction
The rapid growth of web applications and the expansion of cloud-based services have escalated the cybersecurity threat landscape, with WebShells emerging as a significant concern. WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018). According to a Cisco Talos report (Talos, 2024), threat groups used a variety of web shells against vulnerable or unpatched web applications in 35% of incidents in Q4 2024, a sharp increase from the previous quarter, when such activity was observed in only 10% of incidents. These attacks are particularly dangerous due to the stealthy nature of WebShells, which are rapidly evolving to evade traditional detection methods (Hannousse & Yahiouche, 2021).
@@
In response to this challenge, the community has proposed several approaches. Rule-based methods that rely on predefined signatures or heuristics are increasingly ineffective against the complexity and diversity of modern WebShells (Le et al., 2021; Jinping et al., 2020). Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats. However, they require extensive training on large datasets - resources that are often difficult to obtain and sensitive in nature  (Shang et al., 2024). In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks  (Jinping et al., 2020; Zhang et al., 2025).
@@
Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation  (Ma et al., 2024) and vulnerability detection  (Liu & He, 2023; Wang et al., 2025). Studies have shown that with timely engineering, LLMs can perform remarkably well without additional training  (Nong et al., 2024; Trad & Chehab, 2025). They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024). Despite these advantages, there has been limited exploration of LLMs for WebShell detection.
@@
Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks. WebShells often employ obfuscation and encryption techniques, and are embedded in large codebases dominated by benign content (Liu & He, 2023). For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024). In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicates the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024). While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
@@
First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art machine learning (SOTA) methods. Specifically, we evaluate seven closed-source and open-source LLMs of different sizes, including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024) on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples). Our analysis reveals several key findings:
  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near perfect precision (close to 100%), but struggle with recall (e.g., GPT-4’s recall is 85.98%), indicating difficulty in detecting all malicious instances.
  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, have high recall (close to 100%) but suffer from low precision (e.g. Qwen 2.5 Coder 3B’s precision is 38.93%), indicating a tendency to misclassify benign files as malicious.
@@
We compare the performance of LLMs with traditional methods, including Glove+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023). The best performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%. [^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts. In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results..
Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and to improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall. Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors. Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.32%, respectively. In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive or even superior to traditional methods.
@@ 2.2 Context-Aware Code Extraction
7:        Extract context window $[p-\tau,p+\tau]$ from $\mathcal{C}$
@@ 3.2 Experiment Setup
We randomly selected 60% of the dataset to create a fixed known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing webshells from benign programs. These scores, detailed in the Table 3 in the Appendix C, guided the selection of ICL demonstrations from the known sample library.
@@ 4 Results and Analysis
In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection, and assess the improvements provided by our proposed BFAD framework. Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
  • RQ1: How do large and small scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
@@ 4.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
We evaluated seven LLMs, including large-scale models such as GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B, and small-scale models such as Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, and Qwen-2.5-0.5B. These were compared with traditional ML and DL baselines, including sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT). The results are detailed in Table 4 in Appendix D.
Characteristics of LLMs in WebShell Detection
Our evaluation reveals distinct performance characteristics of LLMs in WebShell detection that are influenced by their size. Compared to traditional machine learning and deep learning methods, especially GAT, vanilla LLMs exhibit unbalanced performance - large models lack sufficient recall, while small models lack precision - making them less effective without targeted improvements.
In particular, large LLMs such as GPT-4 and Qwen-2.5-Coder-14B achieve high precision (100% and 99.32%, respectively) but moderate recall (85.98% and 93.63%, respectively), reflecting a conservative classification bias. This is likely due to their extensive training on diverse datasets, which prioritizes accurate identification of benign code over detection of rare or novel WebShells. Code-specific fine-tuning improves the performance of Qwen-2.5-Coder-14B over general-purpose models such as LLaMA-3.1-70B (97.31% precision, 92.36% recall), underscoring the benefits of domain specialization. In contrast, small-scale LLMs such as Qwen-2.5-0.5B and Qwen-2.5-1.5B exhibit nearly perfect recall (100% and 95.77%, respectively) but extremely low precision (18.65% and 34.61%, respectively). This imbalance may be due to their increased sensitivity to prompts, making them more likely to provide answers that match user expectations, a problem exacerbated by their limited ability to model complex code relationships.
Effects of BFAD on LLMs
The BFAD framework significantly improves the performance of LLMs for WebShell detection. For large-scale LLMs, BFAD increases recall while maintaining near-perfect precision; in particular, GPT-4’s recall increases by 12.73% to 98.71%, yielding an F1 score of 99.35%, which exceeds the graph-based GAT baseline (98.87% F1), while LLaMA-3.1-70B achieves an F1 score of 98.40%, closely matching GAT. Conversely, for small LLMs, BFAD significantly increases precision without compromising high recall; Qwen-2.5-0.5B’s precision increases by 52.36% to 71.10%, resulting in an F1 score of 82.67% (a 51.23% improvement), and Qwen-2.5-1.5B achieves an F1 score of 65.33% (up 14.85%). These advances enable BFAD-enhanced LLMs to match or exceed traditional methods without the need for additional fine-tuning or training, driven by a strategic focus on critical code regions and the use of weighted behavioral profiling to select contextually relevant demonstrations. This approach effectively mitigates the inherent context length constraints of LLMs and improves their generalization from limited examples. Notably, small-scale LLMs exhibit more pronounced gains, underscoring the critical role of structured input for models with limited capacity, while large-scale LLMs benefit from improved recall, making them suitable for safety-critical applications. These results highlight the transformative potential of integrating domain-specific strategies with LLMs to address specialized challenges, such as WebShell detection, with exceptional effectiveness.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy. Five demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
@@
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 23.54% reduction from the no-ICL baseline of 84.37%. This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision (100.00%), but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.

### v2

@@ Abstract
WebShell attacks, where malicious scripts are injected into web servers, pose a significant cybersecurity threat. Traditional machine learning and deep learning methods are often hampered by challenges such as the need for extensive training data, catastrophic forgetting, and poor generalization. Recently, Large Language Models (LLMs) have emerged as a powerful alternative for code-related tasks, but their potential in WebShell detection remains underexplored. In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain. Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles. Our results show that, stemming from their distinct analytical strategies, larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off. However, all baseline models lag behind previous State-Of-The-Art (SOTA) methods. With the application of BFAD, the performance of all LLMs improves significantly, yielding an average F1 score increase of 13.82%. Notably, larger models like GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B now outperform SOTA benchmarks, while smaller models such as Qwen-2.5-Coder-3B achieve performance competitive with traditional methods. This work is the first to explore the feasibility and limitations of LLMs for WebShell detection and provides solutions to address the challenges in this task.
@@ 1 Introduction
The rapid growth of web applications and the expansion of cloud-based services have intensified the cybersecurity threat landscape, with WebShells emerging as a significant concern. WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018). According to a Cisco Talos report (Talos, 2024), threat groups used a variety of web shells against vulnerable or unpatched web applications in 35% of incidents in Q4 2024, a sharp increase from the previous quarter, when such activity was observed in only 10% of incidents. These attacks are particularly dangerous due to the stealthy nature of WebShells, which are rapidly evolving to evade traditional detection methods (Hannousse & Yahiouche, 2021).
@@
In response to this challenge, the community has proposed several approaches. Rule-based methods that rely on predefined signatures or heuristics are increasingly ineffective against the complexity and diversity of modern WebShells (Le et al., 2021; Jinping et al., 2020). Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats. However, they require extensive training on large datasets–resources that are often difficult to obtain and sensitive in nature (Shang et al., 2024). In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks (Jinping et al., 2020; Zhang et al., 2025).
@@
Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation (Ma et al., 2024) and vulnerability detection (Liu & He, 2023; Wang et al., 2025). Studies have shown that with effective prompt engineering, LLMs can perform remarkably well without additional training (Nong et al., 2024; Trad & Chehab, 2025). They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024). Despite these advantages, there has been limited exploration of LLMs for WebShell detection.
@@
Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks. WebShells often employ obfuscation and encryption techniques and are embedded in large codebases dominated by benign content (Liu & He, 2023). For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024). In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicate the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024). While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
@@
First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art (SOTA) machine learning methods. Specifically, we evaluate seven closed-source and open-source LLMs of different sizes–including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)–on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples). Our analysis reveals several key findings:
  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near-perfect precision but struggle with recall (e.g., GPT-4’s recall is 85.98%). Their deep contextual analysis, while powerful, can be deceived by well-disguised malicious code, leading to false negatives.
  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, exhibit high recall but suffer from low precision (e.g., Qwen 2.5 Coder 3B’s precision is 38.93%). This stems from their reliance on surface-level pattern matching, which often flags any use of sensitive functions without sufficient contextual validation.
@@
For this comparison, we benchmark against several traditional methods, including GloVe+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023). The best-performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%. [^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts. In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results.
Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall. Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors. Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.23%, respectively. In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive with or even superior to traditional methods.
@@ 2.2 Context-Aware Code Extraction
7:   Extract context window $p-\tau,p+\tau]$ from $\mathcal{C}$
@@ 3.2 Experiment Setup
We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs. These scores, detailed in the Table 4 in the Appendix C, guided the selection of ICL demonstrations from the known sample library.
@@ 4 Results and Analysis
In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection and assess the improvements provided by our proposed BFAD framework. Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
  • RQ1: How do large and small-scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
@@ 4.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
To answer our first research question, we evaluated seven LLMs against traditional ML and DL baselines. The LLMs included large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and small-scale models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B). Baselines consisted of sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT). The detailed performance comparison is presented in Table 1.
Table 1: Performance Comparison of BFAD-Enhanced Models Against Baselines.
Category Model Accuracy Precision Recall F1 Score
Sequence Baselines GloVe+SVM 96.20% 93.30% 94.30% 93.80%
CodeBERT+RF 96.30% 94.00% 95.60% 94.80%
Graph Baselines GCN 96.90% 94.40% 95.30% 94.90%
GAT 98.37% 99.52% 97.39% 98.87%
LLM Baselines (Large) GPT-4 97.27% 100.00% 85.98% 92.46%
LLaMA-3.1-70B 98.01% 97.31% 92.36% 94.77%
Qwen-2.5-Coder-14B 98.64% 99.32% 93.63% 96.39%
Qwen-2.5-3B 93.72% 78.03% 91.84% 84.37%
Qwen-2.5-1.5B 43.62% 34.61% 95.77% 50.84%
Qwen-2.5-0.5B 19.47% 18.65% 100.00% 31.44%
LLM + BFAD GPT-4 99.75% 100.00% 98.71% 99.35% (+6.89)
LLaMA-3.1-70B 99.38% 98.72% 98.09% 98.40% (+3.63)
Qwen-2.5-Coder-14B 98.76% 98.68% 94.90% 96.75% (+0.36)
Qwen-2.5-Coder-3B 78.89% 46.67% 100.00% 63.64% (+7.71)
Qwen-2.5-3B 97.39% 88.64% 99.36% 93.69% (+9.32)
Qwen-2.5-1.5B 80.40% 48.51% 100.00% 65.33% (+14.49)
Qwen-2.5-0.5B 91.94% 71.10% 98.73% 82.67% (+51.23)
Performance Summary
Our experimental results yield two core findings. First, vanilla LLMs exhibit a stark, size-dependent trade-off between precision and recall, with all models lagging behind the state-of-the-art GAT baseline. For instance, GPT-4 achieves 100% precision but only 85.98% recall, while the small Qwen-2.5-0.5B model reaches 100% recall but a mere 18.65% precision. This indicates that out-of-the-box LLMs are not well-calibrated for the nuanced task of WebShell detection, making targeted intervention necessary for effective deployment.
Second, the BFAD framework dramatically improves performance across all models, elevating large LLMs to surpass the GAT baseline and making small models viable competitors. With BFAD, GPT-4’s F1 score increases by 6.89 points to 99.35%, exceeding the GAT baseline. The most significant gain is seen in Qwen-2.5-0.5B, whose F1 score improves by a remarkable 51.23 points. This demonstrates that by strategically guiding the model’s focus on behaviorally relevant code, BFAD effectively bridges the performance gap, unlocking the potential of LLMs for this task without resource-intensive retraining.
Analysis of Model Behavior: Misclassification Patterns
The performance disparities observed in our results appear to stem from fundamental differences in the models’ analytical reasoning. The precision-recall imbalance in baseline LLMs is a direct reflection of their reasoning strategy. Small models like Qwen-2.5-0.5B tend to rely on surface-level heuristics; they identify keywords like base64_decode and flag them as malicious without assessing the surrounding code’s intent. This leads to high recall but generates numerous false positives when such functions are used legitimately (e.g., in a cryptography library). In contrast, large models like LLaMA-3.1-70B employ deep contextual reasoning, correctly identifying legitimate use cases and thus achieving high precision. However, this sophistication can be a double-edged sword, as these models can be misled by well-disguised WebShells, rationalizing away malicious indicators and causing false negatives that reduce recall. Our proposed BFAD mitigates these issues by guiding smaller models to better understand function-level behaviors and helping larger models focus on critical functions. As a result, it significantly improves detection performance across scales.
Analysis of Model Behavior: Counterintuitive Scaling
We observed a counterintuitive scaling effect where the Qwen-2.5-1.5B model, despite outperforming the 0.5B variant in baseline tests, performed worse with BFAD. We attribute this to differing sensitivities to ICL across scales. The smallest models, like the 0.5B variant, are highly susceptible to label bias and tend to mimic the ICL demonstration’s label without deep reasoning (Fei et al., 2023). Because BFAD provides accurately matched demonstrations via WBFP, the 0.5B model benefits disproportionately from this behavior. In contrast, the 1.5B model begins to exhibit more independent reasoning based on its pre-training knowledge. This emergent but still imperfect reasoning can conflict with the ICL example’s guidance, degrading performance. This aligns with prior findings that LLMs sometimes rely more on pre-training than on provided examples (Peng et al., 2025). Therefore, the 0.5B model’s superior performance with BFAD does not indicate greater intelligence but is an artifact of its strong alignment with the high-quality examples provided.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
@@
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 23.54% reduction from the no-ICL baseline of 84.37%. This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 16.24% decrease from the No-ICL baseline of 92.46%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 85.98%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.